### v1

@@ 3.2 Convolutional Neural Networks
In the introduction of this paper, we established the potential of NNs for predicting $\rm R_{ta}$, we now focus on designing a model specifically tailored to the nature of our data. The observational proxies for $\rm R_{ta}$ are inherently spatial in nature, represented as 2D projections of galaxy distributions, along with their masses and their velocities.
This spatial structure necessitates a NN architecture capable of effectively capturing and analyzing such features. In conventional NNs, the layers are fully connected, designed primarily for one-dimensional inputs. In this work, we use Convolutional Neural Networks (CNNs), which are best suited for single and multi-channel images (2D and 3D inputs). CNNs use convolutional layers which connect the neurons of one layer only to a small region of the neurons of the previous layer (e.g. pixels of the image), allowing for the recognition of low level spacial features. The main structure of our NN can be seen in Table 2. This structure was the best candidate in almost all of the trials performed as it was favored by our optimization study.
@@ 3.3 Metrics
To evaluate the accuracy of the predicted $\rm R_{ta}$, we utilized established statistical metrics: the coefficient of determination (R2 score) and Shapley values.
We use the $\rm R^{2}$ score to evaluate the accuracy of the model, by comparing the predicted and the true $\rm R_{ta}$. An R2 score of 1.0 signifies that the model accurately predicts the target variable ($\rm R_{ta}$) based on the feature variables (projections). Conversely, an R2 score of 0 indicates that the model does not predict the target variable at all, i.e., the model is no better than simply taking the mean of the target variable distribution. Notably, the R2 score can become negative when the chosen model fits the data worse than a horizontal hyperplane.
In the context of machine learning, and the interpretability of its models, Shapley values provide a method to allocate the contribution of each feature to the prediction for each individual instance. The Shapley value of a feature can be interpreted as its average marginal contribution to the prediction, across all possible combinations of features.
@@ 4 Results
In this section, we detail the outcomes of our analysis. Our primary focus is assessing the predictive capability of our trained CNN in determining the $\rm R_{ta}$ of simulated clusters, utilizing their various mock projection images. We specifically examine the network’s effectiveness, trained on different mock data, including the mass distribution of galaxies proximal to the clusters, as well as their velocity distribution, the surrounding galactic number density, and varying levels of halo projection around the clusters as viewed in the plane of the sky (initially the 3 types of histograms of Sect. 3.1 with the 2 velocity cuts, while considering the $\rm 4^{th}$ type in Sect. 4.3). Additionally, we consider the impact of the central cluster region data on our model and, we evaluate the model’s predictive accuracy when tested in a cosmological setting different from the one in which it is trained.
Refer to captionRefer to caption
Figure 2: Performance comparison across different models. Upper panel: Comparison between low- and high- velocity-cut models performance. The plot shows different $\rm R^{2}$ values for each combination of the 3 features considered: mass column density, number column density and mean line-of-sight velocity (indicated as “mass”, “num”, and “vel” respectively). Best results are obtained for the low velocity cut (due to less contamination) and for data containing the mass. Lower panel: Comparison between models including or excluding the central halo. Both analyses are performed for the low-velocity-cut data. Without the central halo, performance drops significantly across models; all models without it predict the turnaround with similar score.
Refer to captionRefer to captionRefer to caption
Figure 3: Comparison between predicted and true values of $\rm R_{ta}$ for different CNN models, when using mass information. The colors represent the number density of the plotted points, calculated using a Gaussian kernel density estimate. While the models in the left and middle panel show some correlation between true predicted $\rm R_{ta}$ values, the right-panel model appears to mostly predict the mean value of the $\rm R_{ta}$ distribution.
@@ 4.1 General
Initially, we focused on analyzing projected cluster images from the MDPL2 simulation at redshift $z=0$, applying both low and high velocity cuts to get a general idea of the potential of our methodologies. As outlined in Sect. 3.2, CNNs can be utilized with both single and multi-channel inputs, enabling us to merge our 3 types of post-processed single-channel mock data images (mass column density, number column density, mean line-of-sight velocity) in various configurations to generate multi-channel images. For instance, combining all three data types can be thought of as an RGB image, with the R, G, and B channels representing our distinct data types.
The outcomes of training the network with all possible data combinations are illustrated in the upper panel of Fig. 2, depicted as a histogram of the $\rm R^{2}$ scores derived from the test set for each combination, displayed along the x-axis. Training with the low velocity cut data yielded the highest accuracy, as evidenced by the superior $\rm R^{2}$ scores marked by blue bars. This result was anticipated, given the increased projection depth contamination associated with the high velocity cut.
Models incorporating mass column density images demonstrated the highest accuracy, as indicated by the first four bar pairs, aligning with expectations that the gravitational influence of a cluster’s mass correlates with the distance from the cluster’s center where the Hubble flow becomes dominant. Mean velocity data did not exhibit a strong correlation with the $\rm R_{ta}$. We aim to delve deeper into these observations in Sects. 4.2 and 4.3, examining the cases of mass column density and velocity in further detail. Moving forward, we will exclusively consider the low-velocity-cut data, corresponding to the high-accuracy-redshift case, aiming to examine the most favorable scenario possible.
@@ 4.2 Mass distribution
In the previous section, a significantly better performance of the models trained on data containing the mass was established. This leads us to investigate what the model’s performance would be without information on the central, more massive halo (i.e. without information on the mass of the central collapsed galaxy cluster).
@@
To this end, we removed the central halo completely from all of the regions and re-trained our model with the low-velocity-cut data from the MDPL2 $\rm z=0$ output. The results can be seen in the lower panel of Fig. 2. Without the central halo, the performance of the mass-containing models drop significantly: there is a very noticeable difference between the height of the blue and red bars in the first 4 models. This implies that their better performance was due to the information contained in the central halo’s mass. All of the other models predict the turnaround with about the same score.
@@
A more accurate representation of these results is shown in Fig. 3 with scatter plots between the predicted values of the turnaround radii and their real values. If the model was perfect then all the points ($\rm R_{\rm ta,predicted}$, $\rm R_{\rm ta,true}$) would be located on the $\rm y=x$ (black solid) line. In the case where the central halo is missing (right panel) there seems to be a little to no correlation whatsoever. This result underscores that what the model is extracting and learning in the first two cases (left and middle panel) is a correlation between the central mass and the turnaround radius (which has been shown to exist analytically, Korkidis & Pavlidou 2024) rather than some actual image feature at the location of the turnaround radius itself.
@@
We also evaluated our model in cosmological settings different than the one our model was trained. The $\rm R^{2}$ score from different redshifts from MDPL2 and from different cosmologies from both MDPL2 and the Virgo Simulations can be seen in the upper panel of Fig. 4 (left and right panel respectively) where larger circles correspond to higher $\rm R^{2}$ scores. Best models seem to be the ones trained with the data from MDPL2, most likely due to the larger training dataset. Additionally, a model trained at one redshift is not successful when tested at a different one. Models perform worse as the redshift of the test set moves further away from the redshift of the training set, as indicated by the decreasing size of the circles as we move away from the diagonal in the upper left panel of Fig. 4. If we accept that the model is primarily reconstructing the scaling of the central massive halo with the turnaround radius, this behavior is expected because of the redshift dependence of $\rm\rho_{ta}$, and of its scaling with the central halo mass.
Refer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 4: $\rm R^{2}$ scores of different combinations of training and testing data using the mass column density (upper panel) or velocity dispersion (lower panel) information (size of circles indicates $\rm|R^{2}|$ for easier comprehension of general trends). 80% of the simulation data were used for training-validation and 20% was used for the testing in each case located in the diagonal. Left panel: Training and testing data from different MDPL2 redshifts (z=0, 0.49, 1.03). As expected, for each training set the $\rm R^{2}$ score is the highest when evaluated on data from the same redshift. Since MDPL2 is a $\rm\Lambda CDM$ concordance cosmological simulation, the $\rm\rho_{ta}$ depends on redshift. Therefore, for different redshifts the model cannot find an apparent correlation between $\rm M_{vir}$ and $\rm R_{ta}$ because $\rm\rho_{ta}$ varies with redshift. The velocity information plot has the same form as the mass, indicating that the model just correlates the mass with the velocity dispersion. Right panel: Training and testing data from different simulations and cosmological models (MDPL2, Virgo $\rm\Lambda CDM$, Virgo OCDM, Virgo SCDM and the merged data from MDPL2 and Virgo $\rm\Lambda CDM$). Best models seem to be the ones containing the data from the MDPL2 possibly due to the fact of the larger number of training instances.
Refer to captionRefer to captionRefer to caption
Figure 5: Comparison between predicted and true values of of the $\rm R_{ta}$ for models using the ”stacked” images from 3000 bins of $\rm R_{ta}$ from: the MDPL2 data (left panel); the merged MDPL2 with Virgo $\rm\Lambda CDM$ data (middle panel); the merged MDPL2 with Virgo $\rm\Lambda CDM$ data with removed information inside the $\rm R_{200}$ of the central overdensity (right panel). $\rm R^{2}$ scores are 0.36, 0.57 and 0.59 respectively. The colors represent the number density of the plotted points, calculated using a Gaussian kernel density estimate. It is apparent that the merging of the 2 datasets significantly improves the performance of the model even without the central halos’ velocity information.
@@ 4.3 Velocity distribution
In Sect. 4.1, we observed that velocity data alone did not yield significant information. A model trained and tested solely on velocity data performed unsatisfactorily. Given the known correlation between the mass of the central halo and $\rm R_{ta}$, and also between velocity dispersion and mass of the central halo, one would rather anticipate a correlation between the velocity dispersion of surrounding halos and the central halo’s mass, and consequently, the turnaround radius. To address this, we explore a $\rm 4^{th}$ type of data: the velocity dispersion.
@@ 4.3.1 Stacking
Our initial focus was on the MDPL2 data at $\rm z=0$. We hypothesized that the sparsity of galaxies in computing a velocity dispersion image might obscure any relationship with the $\rm R_{ta}$. To counteract this, we employed a “stacking” technique for our projections, based on their respective $\rm R_{ta}$ values. This stacking involved categorizing the $\rm R_{ta}$ into bins and merging data from clusters within the same bin. We then assigned a new $\rm R_{ta}$ value for each bin, calculated as the mean of the $\rm R_{ta}$ values within it. The motivation here is that clusters normalized with respect to overdensity radii or masses are known to exhibit similar dynamical characteristics. Subsequently, as detailed in Sect. 3.1, we transformed these data into images. This time, however, we focused on computing the standard deviation of the velocities within each bin.
After incorporating the new data type into our training process, we reassessed the model’s performance on the test set. The left panel of Fig. 5 displays the comparison between the true values and those predicted by the model, which was trained using 3000 $\rm R_{ta}$ bins. This approach yielded approximately 1200 cluster images, as some bins were devoid of data. Our analysis revealed a modest improvement in correlation, as evidenced by the increase in the correlation coefficient (from $\rm R^{2}=0.24$ to $\rm R^{2}=0.36$, comparing with the one of Fig. 2). Although this enhancement in correlation is not markedly significant, it suggests that a stacking approach, based on variables influencing cluster dynamics, holds promise for future research applications. [^1] The results of a similar approach but by stacking images based on their $\rm R_{200}$ can be found in Fig. 9 of the appendix.

### v2

@@ 3.2 Convolutional neural networks
In the introduction of this paper, we established the potential of NNs for predicting $\rm R_{ta}$. We now focus on designing a model specifically tailored to the nature of our data. The observational proxies for $\rm R_{ta}$ are inherently spatial in nature, represented as 2D projections of galaxy distributions, along with their masses and their velocities.
This spatial structure necessitates a NN architecture capable of effectively capturing and analyzing such features. In conventional NNs, the layers are fully connected, designed primarily for one-dimensional inputs. In this work, we used convolutional neural networks (CNNs), which are best suited for single and multi-channel images (2D and 3D inputs). CNNs use convolutional layers, which connect the neurons of one layer only to a small region of the neurons of the previous layer (e.g., pixels of the image), allowing for the recognition of low-level spatial features. Although a simple NN with fully connected layers (as a universal function approximator) would have been able to approximate any function, we expect spatial correlations in our data (since halos trace the underlying dark matter density field), which are captured more easily with CNNs.
Graph neural networks (GNNs) are a promising alternative. They operate on graphs, which consist of nodes representing entities (e.g., in our case, halos characterized by positions, masses, and velocities), edges encoding relationships between these entities, and potentially global context features (e.g., cosmological parameters or background overdensity). The graph’s connectivity, defined by the edges, plays a crucial role in the model’s performance. In our case, the most natural and widely adopted approach would be to construct this graph based on the proximity of the halos. Two commonly used and effective choices are a k-nearest neighbors (k-NN) graph since it ensures connectivity between the nodes and a linking radius approach, as shown to sometimes produce better results for similar applications (e.g., Villanueva-Domingo & Villaescusa-Navarro 2022). Preliminary tests using GNNs with graphs constructed with both approaches (for $\rm k\in\{5,10\}$ and a linking radius $\rm\in\{1,2,3\}\;Mpc$) did not show any improvement over CNNs.  [^4] We used PyTorch Geometric (Fey & Lenssen 2019) on the MDPL2 dataset, using the halo mass and/or line-of-sight velocity as node features. The best GNN model consisted of two graph convolutional network (GCN) layers with 64 and 16 neurons, respectively, each followed by ReLU activation and dropout (30 %), followed by a global mean pooling operation and a fully connected layer of eight neurons. The network was trained using the Adam optimizer (Kingma & Ba 2014) with a learning rate of 0.1.
The main structure of our CNN can be seen in Table 2. This structure was the best candidate in almost all of the trials performed, as it was favored by our optimization study.
@@ 3.3 Metrics
To evaluate the accuracy of the predicted $\rm R_{ta}$, we utilized established statistical metrics: the coefficient of determination (R2 score) and Shapley values (Shapley 1953). Both are explained in more detail in the following paragraphs.
We used the $\rm R^{2}$ score to evaluate the accuracy of the model, by comparing the predicted and the true $\rm R_{ta}$, it is defined as
$\rm R^{2}=1-\frac{\sum_{i}(R^{true}_{ta,i}-R^{pred}_{ta,i})^{2}}{\sum_{i}(R^{true}_{ta,i}-\langle R^{true}_{ta,i}\rangle)^{2}},$ (3)
where i are samples from the test set. An R2 score of 1 signifies that the model accurately predicts the target variable ($\rm R_{ta}$) based on the feature variables (projections). Conversely, an R2 score of 0 indicates that the model does not predict the target variable at all; in other words, the model is no better than simply taking the mean of the target variable distribution. Notably, the R2 score can become negative for arbitrarily worse models when the chosen model not only fails to predict the truth, but even predicts a number further away from the mean.
In the context of machine learning and the interpretability of its models, Shapley values provide a method to allocate the contribution of each feature (pixels in our case) to the prediction for each individual instance. In this work, we used the SHAP library (Lundberg & Lee 2017), which provides an efficient implementation of Shapley value estimation. The main idea is to approximate the prediction of a complex black-box model $\rm f$ (in our case, a NN), with a simpler interpretable explanation model $\rm g$. SHAP uses a simplified vector of binary input variables $\rm x^{\prime}\in\{0,1\}^{M}$, indicating the presence or absence of each of the $\rm M$ inputs. These are mapped to the real input $\rm x$ through a function $\rm x=h_{x}(x^{\prime})$, which fills in missing features using their expected distribution. The explanation model is defined as
$\rm g(z^{\prime})=\phi_{0}+\sum_{i=1}^{M}\phi_{i}z^{\prime}_{i},$ (4)
where $\rm z^{\prime}$ is a binary vector (similar to $\rm x^{\prime}$), $\rm g(z^{\prime})\approx f(h_{x}(z^{\prime}))\text{ when }z^{\prime}\approx x^{\prime}$, and $\phi_{0}=\mathbb{E}[f(x)]$ is the expectation value of the model, given a specific dataset. The only solution to the values $\phi_{i}$ that satisfy the desirable properties (described in Lundberg & Lee 2017) are the Shapley values:
$\rm\phi_{i}(f,x)=\sum_{z^{\prime}\subseteq x^{\prime}}\frac{|z^{\prime}|!\,(M-|z^{\prime}|-1)!}{M!}\left[f(h_{x}(z^{\prime}_{+i}))-f(h_{x}(z^{\prime}_{-i}))\right],$ (5)
where $\rm|z^{\prime}|$ is the number of non-zero elements in $\rm z^{\prime}$, $\rm z^{\prime}\subseteq x^{\prime}$ indicates the sum over all possible $\rm z^{\prime}$ that have a lower or equal number of non-zero elements to $\rm x^{\prime}$, and $\rm z^{\prime}_{+i}$ ($\rm z^{\prime}_{-i}$) is the input vector including (excluding) feature $\rm i$.
In practice, SHAP quantifies how much each input feature pushes the model’s prediction above or below the expected value $\rm\phi_{0}$. For this study, we used the DeepExplainer, a gradient-based variant of the SHAP library, to efficiently approximate Shapley values.
@@ 4 Results
In this section, we detail the outcomes of our analysis. Our primary focus was to assess the predictive capability of our trained CNN in determining the $\rm R_{ta}$ of simulated clusters, utilizing their various mock projection images. We specifically examined the network’s effectiveness, trained on different mock data, including the mass distribution of galaxies in the cluster regions, as well as their velocity distribution, the surrounding galactic number density, and varying levels of halo projection around the clusters as viewed in the plane of the sky (initially the three types of histograms of Sect. 3.1 with the two velocity cuts, while considering the fourth type in Sect. 4.3). Additionally, we considered the impact of the central cluster region data on our model, and we evaluated the model’s predictive accuracy when tested in a cosmological setting different from the one in which it was trained.
@@ 4.1 General
Initially, we focused on analyzing projected cluster images from the MDPL2 simulation at redshift $z=0$, applying both low and high velocity cuts to get a general idea of the potential of our methodologies. As outlined in Sect. 3.2, CNNs can be utilized with both single and multi-channel inputs, enabling us to merge our three types of post-processed single-channel mock data images (mass column density, number column density, mean line-of-sight velocity) in various configurations to generate multi-channel images. For instance, combining all three data types can be thought of as an RGB image, with the R, G, and B channels representing our distinct data types.
The outcomes of training the network with all possible data combinations are illustrated in the upper panel of Fig. 2, depicted as a histogram of the $\rm R^{2}$ scores derived from the test set for each combination, displayed along the x axis. Training with the low-velocity-cut data yielded the highest accuracy, as evidenced by the superior $\rm R^{2}$ scores marked by blue bars. This result was anticipated, given the increased projection depth contamination associated with the high velocity cut.
Models incorporating mass column density images demonstrated the highest accuracy, as indicated by the first four bar pairs, aligning with expectations that the gravitational influence of a cluster’s mass correlates with the distance from the cluster’s center, where the Hubble flow becomes dominant. Mean velocity data did not exhibit a strong correlation with the $\rm R_{ta}$. We delve deeper into these observations in Sects. 4.2 and 4.3, examining the cases of mass column density and velocity in further detail. Moving forward, we exclusively consider the low-velocity-cut data, corresponding to the high-accuracy-redshift case, aiming to examine the most favorable scenario possible.
@@ 4.2 Mass distribution
In the previous section, a significantly better performance of the models trained on data containing the mass was established. This led us to investigate what the model’s performance would be without information on the central, more massive halo (i.e., without information on the mass of the central collapsed galaxy cluster).
@@
To this end, we removed the central halo completely from all of the regions and re-trained our model with the low-velocity-cut data from the MDPL2 $z=0$ output. The results can be seen in the lower panel of Fig. 2. Without the central halo, the performance of the mass-containing models drops significantly: there is a very noticeable difference between the height of the blue and red bars in the first four models. This implies that their better performance was due to the information contained in the central halo’s mass. All of the other models predict the turnaround with about the same score.
@@
A more accurate representation of these results is shown in Fig. 3 with scatter plots between the predicted values of the turnaround radii and their real values. If the model was perfect then all the points ($\rm R_{\rm ta,predicted}$, $\rm R_{\rm ta,true}$) would be located on the $\rm y=x$ (solid black) line. In the case where the central halo is missing (right panel), there seems to be little to no correlation whatsoever. This result underscores that what the model is extracting and learning in the first two cases (left and middle panel) is a correlation between the central mass and the turnaround radius (which has been shown to exist analytically, Korkidis & Pavlidou 2024) rather than some actual image feature at the location of the turnaround radius itself.
Figure 2: Performance comparison across different models. Upper panel: Comparison between low- and high-velocity-cut models’ performance. The plot shows different $\rm R^{2}$ values for each combination of the three features considered: mass column density, number column density, and mean line-of-sight velocity (indicated as “mass,” “num,” and “vel” respectively). Best results are obtained for the low velocity cut (due to less contamination) and for data containing the mass. Lower panel: Comparison between models including or excluding the central halo. Both analyses are performed for the low-velocity-cut data. Without the central halo, performance drops significantly across models; all models without it predict the turnaround with a similar score.
Figure 3: Comparison between predicted and true values of $\rm R_{ta}$ for models using the MDPL2 “mass” images from: the low-velocity-cut data (left panel); the high-velocity-cut data (middle panel); and the low-velocity-cut data after removing the central halo of each projection (right panel). The colors represent the number density of the plotted points, calculated using a Gaussian kernel density estimate. $\rm R^{2}$ scores and the mean absolute errors (MAEs) are shown at the bottom right of each plot. While the models in the left and middle panel show some correlation between true and predicted $\rm R_{ta}$ values, the right-panel model appears to mostly predict the mean value of the $\rm R_{ta}$ distribution, especially when comparing the MAEs with the standard deviation of the $\rm R_{ta}$ distribution ($\rm 2.2\;Mpc$) as shown in the upper panel of Fig. 1.
@@
We also evaluated our model in cosmological settings different than the one our model was trained on. The $\rm R^{2}$ score from different redshifts from MDPL2 and different cosmologies from both MDPL2 and the Virgo simulations can be seen in the upper panel of Fig. 4 (left and right panel, respectively). Best models seem to be the ones trained with the data from MDPL2, most likely due to the larger training dataset. Additionally, a model trained at one redshift is not successful when tested at a different one. Models perform worse as the redshift of the test set moves further away from the redshift of the training set, as indicated by the decreasing $\rm R^{2}$ score as we move away from the diagonal in the upper left panel of Fig. 4. If we accept that the model is primarily reconstructing the scaling of the central massive halo with the turnaround radius, this behavior is expected because of the redshift dependence of $\rm\rho_{ta}$, and of its scaling with the central halo mass.
@@
Figure 4: $\rm R^{2}$ scores of different combinations of training and testing data using the mass column density (upper panel) or velocity dispersion (lower panel) information. 80% of the simulation data were used for training-validation, and 20% was used for the testing in each case located in the diagonal. Left panel: Training and testing data from different MDPL2 redshifts (z=0, 0.49, 1.03). As expected, for each training set, the $\rm R^{2}$ score is the highest when evaluated on data from the same redshift. Since MDPL2 is a $\rm\Lambda CDM$ concordance cosmological simulation, the $\rm\rho_{ta}$ depends on redshift. Therefore, for different redshifts, the model cannot find an apparent correlation between $\rm M_{vir}$ and $\rm R_{ta}$ because $\rm\rho_{ta}$ varies with redshift. The velocity information plot has the same form as the mass, indicating that the model just correlates the mass with the velocity dispersion. Right panel: Training and testing data from different simulations and cosmological models (MDPL2, Virgo $\rm\Lambda CDM$, Virgo OCDM, Virgo SCDM, and the merged data from MDPL2 and Virgo $\rm\Lambda CDM$). Best models seem to be the ones containing the data from the MDPL2, possibly due to the fact of the larger number of training instances. Negative $\rm R^{2}$ values are a direct consequence of the fact that evaluating a model on another simulation can predict values even worse than the mean (as discussed in Sect. 3), since the target value distribution is an unknown.
@@ 4.3 Velocity distribution
In Sect. 4.1, we observed that velocity data alone did not yield significant information. A model trained and tested solely on velocity data performed unsatisfactorily. Given the known correlation between the mass of the central halo and $\rm R_{ta}$, and also between velocity dispersion and mass of the central halo, one would rather anticipate a correlation between the velocity dispersion of surrounding halos and the central halo’s mass, and consequently, the turnaround radius. To address this, we explored a fourth type of data: the velocity dispersion.
@@ 4.3.1 Stacking
Our initial focus was on the MDPL2 data at $z=0$. We hypothesized that the sparsity of galaxies in computing a velocity dispersion image might obscure any relationship with the $\rm R_{ta}$. To counteract this, we employed a “stacking” technique for our projections, based on their respective $\rm R_{ta}$ values. This stacking involved categorizing the $\rm R_{ta}$ into bins and merging data from clusters within the same bin. We then assigned a new $\rm R_{ta}$ value for each bin, calculated as the mean of the $\rm R_{ta}$ values within it. The motivation here was that clusters normalized with respect to overdensity radii or masses are known to exhibit similar dynamical characteristics. Subsequently, as detailed in Sect. 3.1, we transformed these data into images. This time, however, we focused on computing the standard deviation of the velocities within each bin.
After incorporating the new data type into our training process, we reassessed the model’s performance on the test set. The left panel of Fig. 5 displays the comparison between the true values and those predicted by the model, which was trained using 3000 $\rm R_{ta}$ bins (with a bin width of $\rm 0.0066\;Mpc$). This approach yielded approximately 1200 cluster images, as some bins were devoid of data. Our analysis revealed a modest improvement in correlation, as evidenced by the increase in the correlation coefficient (from $\rm R^{2}=0.24$ to $\rm R^{2}=0.36$, comparing with the one of Fig. 2). Although this enhancement in correlation is not markedly significant, it suggests that a stacking approach, based on variables influencing cluster dynamics, holds promise for future research applications. [^5] The results of a similar approach, but by stacking images based on their $\rm R_{200}$, can be found in Fig. 10 of the appendix.